OpenStudy (zmudz):

A vector v is called a unit vector if ||v||=1. Let a,b, and c be unit vectors, such that a+b+c = 0. Show that the angle between any two of these vectors is \(120^{\circ}\).

OpenStudy (rsadhvika):



OpenStudy (zarkon):

look at \[a\cdot(a+b+c = 0)\] \[b\cdot(a+b+c = 0)\] \[c\cdot(a+b+c = 0)\]